Find the slope and the y-intercept.
y = 4x – 3
Help

Found 2 solutions by Earlsdon, venugopalramana:
Answer by Earlsdon(6103)

The given equation is already in the "slope-intercept" form: y = mx + b
m is the slope and b is the y-intercept. Compare:
y = 4x - 3 and
y = mx + b
You can see that:
m = 4 So the slope is 4
b = -3 So the y-intercept is -3

Answer by venugopalramana(3286)

Find the slope and the y-intercept.
y = 4x – 3
Help
std.eqn.is
y=mx+c...where m=slope and c=y-intercept..hence by comparison
m=4 and -3=y-intercept